OpenStudy (anonymous):

Convert the degree measure to radians. Give your answer as a rational number multiplied by π. a. 120º b. 36º

OpenStudy (inkyvoyd):

hint 360 degrees=2pi radians

OpenStudy (anonymous):

You can do it like this. \[\frac{120º}{x} = \frac{180}{ \pi}\] Then solve x \[x = \frac{120º \pi }{ 180º} = \frac{2 \pi }{ 3}\] Now, can you try (b)? Replace 120º by 36º and solve x. Then you would find the answer.